What is the number of stories in the tallest house you can build with two 52-card decks? How many cards would it take?
Answer by checkley77(12569)

2*52 card decks=104 cards.
Using 2 cards per story you can build 104/2-52 stories.
Using 3 cards per story you can build 104/3=34 stories.
Using 4 cards per story you can build 104/4=26 stories.